OpenStudy (jtug6):

Would someone mind helping me simplify the expression: sin( tan^-1 (x) ) or sin(arctan x)

OpenStudy (neonumbrella5115):

\[\sin (\tan^{-1} x)\] Is this your equation?

OpenStudy (jtug6):

Yep

OpenStudy (anonymous):

let tan^-1(x) =u

OpenStudy (jtug6):

Ok, and then x = tan (u)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

where u is between -pi/2 and pi/2 right?

OpenStudy (jtug6):

right because of tangent?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so then you have sin(u) between -pi/2 and pi/2

OpenStudy (jtug6):

right

OpenStudy (anonymous):

so then the hypotenuse is gonna be sqrt(x^2+1)

OpenStudy (jtug6):

How so? Are you using a right triangle then applying the pythag theorem based off of x/1 = tan(u)?

OpenStudy (anonymous):

u got it

OpenStudy (jtug6):

is there any other method aside using a right tri or is that just the easiest?

OpenStudy (anonymous):

there are other methods but this the one i know sorry

OpenStudy (jtug6):

It's alright. Thanks anyway.

OpenStudy (anonymous):

np

zepdrix (zepdrix):

\[\large\rm \sin\left[\color{orangered}{\arctan(x)}\right]\]Figure it out jtug? :)

OpenStudy (jtug6):

Yes! So u = arc tan x so x = tan u and we know that x = x/1 so x/1 = tan u meaning that x and 1 are opp/adj to tan which we can then use to find sin(u) BECAUSE u = arc tan x which = our original eq sin [arctan(x)] so sin(u) = opp/hyp which we dont have hyp yet meaning we use pythag theorem because we can use a right triangle which c^2 = x^2 + 1 so c = sqrt x^2+1 which gives us our hyp so we're left with x/[sqrt(x^2+1)] :D

OpenStudy (jtug6):

Thanks for asking btw.

OpenStudy (jtug6):

Calc 2 is beautiful so far. difficult, but beautiful.

zepdrix (zepdrix):

ya fun stuff :))